# Graph Neural Network based Log Anomaly Detection and Explanation

### v1

@@ 3.1. Graph-based Log Anomaly Detection
We can use these definitions to define graph-level anomaly detection:
@@
  • A log group (namely a graph) is considered a quantitative anomaly if the occurrence frequencies of some events in the group are higher or lower than expected from what is commonly observed. For example, if a file is opened (event $A$) twice, it should normally also be closed (event $B$) twice. In other words, the number of event occurrences #$\#A=\#B$ in a normal pattern and an anomaly is detected if #$\#A\neq\#B$.
  • A log group (namely a graph) is considered to contain sequential anomalies if the order of certain events violates the normal order pattern. For instance, a file can be closed only after it has been opened in a normal workflow. In other words, the order of event occurrences $A\rightarrow B$ is considered normal while $B\rightarrow A$ is considered anomalous.
An advantage of graph-based anomaly detection is that it detect these two types of anomalies, but also anomalies reflected in the structure of the graphs. Moreover, no unsupervised log anomaly detection approaches represent event logs as attributed, directed, weighted graphs, which allow for even higher expressiveness than undirected graphs (and thus limiting the information loss resulting from the representation of the log files as graphs).
4. Preliminaries: Digraph Inception Convolutional Nets
To learn node representations for attributed, directed, and edge-weighted graphs, (Tong et al., 2020) proposed Digraph Inception Convolutional Networks (DiGCN).
@@ 4. Preliminaries: Digraph Inception Convolutional Nets
(1) $\textbf{Z}^{(k)}=\begin{cases}\textbf{X}\Theta^{(0)}&\text{$k=0$}\\ \Psi\textbf{X}\Theta^{(1)}&\text{$k=1$}\\ \Phi\textbf{X}\Theta^{(k)}&\text{$k\geq 2$},\end{cases}$
where $\Psi=\frac{1}{2}\left(\Pi^{(1)\frac{1}{2}}\textbf{P}^{(1)}\Pi^{(1)\frac{-1}{2}}+\Pi^{(1)\frac{-1}{2}}\textbf{P}^{(1)T}\Pi^{(1)\frac{1}{2}}\right)$ and $\Phi=\textbf{W}^{(k)\frac{-1}{2}}\textbf{P}^{(k)}\textbf{W}^{(k)\frac{-1}{2}}$. Particularly, $\textbf{Z}^{(k)}\in\mathcal{R}^{|\mathcal{V}|\times f}$ denote the convolved output with $f$ output dimension, and $\Theta^{(0)},\Theta^{(1)},\Theta^{(k)}$ represent the trainable parameter matrices.
@@
(2) $\textbf{P}^{(k)}=\begin{cases}\textbf{I}&\text{$k=0$}\\ \tilde{\textbf{D}}^{-1}\tilde{\textbf{A}}&\text{$k=1$}\\ Ins\left((\textbf{P}^{(1)})^{(k-1)}(\textbf{P}^{(1)T})^{(k-1)}\right)&\text{$k\geq 2$},\end{cases}$
@@
, with $Intersect(\cdot)$ denoting the element-wise intersection of two matrices (see (Tong et al., 2020) for computation details). In addition, $\textbf{W}^{(k)}$ is the diagonalized weight matrix of $\textbf{P}^{(k)}$, and $\Pi^{(1)}$ is the approximate diagonalized eigenvector of $\textbf{P}^{(1)}$. Particularly, the approximate diagonalized eigenvector is calculated based on personalised PageRank (Bahmani et al., 2010), with a parameter $\alpha$ to control the degree of conversion from a digraph to an undirected graph. We omit the details to conserve space, and refer to (Tong et al., 2020) for more details.
Figure 1. The Logs2Graphs pipeline. We use attributed, directed, and weighted graphs for representing the log files with high expressiveness, and integrate representation learning and anomaly detection for accurate anomaly detection. We use off-the-shelf methods for log parsing, log grouping, and graph construction.
@@
(3) $\textbf{Z}=\sigma\left(\Gamma\left(\textbf{Z}^{(0)},\textbf{Z}^{(1)},...,\textbf{Z}^{(k)}\right)\right),$
@@
5. Graph-Based Anomaly Detection for Event Logs
@@
First, after collecting logs from a system, the log parsing step extracts log events and log parameters from raw log messages. Since log parsing is not the primary focus of this article, we use Drain (He et al., 2017) for this task. Drain is a log parsing technique with fixed depth tree, and has been shown to generally outperform its competitors (Zhu et al., 2019). We make the following assumptions on the log files:
@@
5.1. Graph Construction
We next explain how to construct an attributed, directed, and edge-weighted graph given a group of parsed log messages, and illustrate this in Figure 2. Particularly, the motivation of graph construction is to keep everything relevant in log data.
First, we construct nodes to represent the different log events. That is, the number of nodes depends on the number of unique log events that occur in the log group. Second, starting from the first line of log messages in chronological order, we add an directed edge from log event $L_{i}$ to $L_{j}$ and set its edge-weight to $1$ if the next event after $L_{i}$ is $L_{j}$. If the corresponding edge already exists, we increase its edge-weight by $1$. In this manner, we obtain a labelled, directed, and edge-weighted graph.
Figure 2. The construction of an attributed, directed, and edge-weighted graph from a group of log messages.
@@ 5.1. Graph Construction
  2. (2)
    Word embedding: we use Glove (Pennington et al., 2014), a pre-trained word embedding model with 200 embedding dimensions to generate a vector representation for each English word in a log event;
  3. (3)
    Sentence embedding: we generate a vector representation for each log event. Considering that the words in a sentence are usually not of equal importance, we use Term Frequency-Inverse Document frequency (TF-IDF) (Ramos et al., 2003) to measure the importance of words. As a result, the weighted sum of word embedding vectors composes the vector representation of a log event.
@@
5.2. OCDiGCN: One-Class Digraph Inception Convolutional Nets
We next describe One-Class Digraph Inception Convolutional Networks, abbreviated as OCDiGCN, a novel method for end-to-end graph-level anomaly detection. We chose to build on Digraph Inception Convolutional Networks (DiGCN) (Tong et al., 2020) for their capability to handle directed graphs, which we argued previously is an advantage in graph-based log anomaly detection.
Considering that DiGCN was designed for node representation learning, we repurpose it for graph representation learning as follows:
(4) $\textbf{z}=\mathrm{Readout}(\textbf{Z}_{i}\mid i\in\{1,2,...,|\mathcal{V}|\}).$
That is, at the final iteration layer, we utilise a so-called $\mathrm{Readout}(\cdot)$ function to aggregate node vector representations to obtain a graph vector representation. Importantly, $\mathrm{Readout}(\cdot)$ can be a simple permutation-invariant function such as maximum, sum or mean, or a more advanced graph-level pooling function (Ying et al., 2018).
Next, note that DiGCN work did not explicitly enable learning edge features (i.e., Y). However, as DiGCN follows the Message Passing Neural Network (MPNN) framework (Gilmer et al., 2017), incorporating Y into Equation (1) and conducting computations in Equations (2-4) analogously enables learning edge features.
@@ 5.2. OCDiGCN: One-Class Digraph Inception Convolutional Nets
In graph anomaly detection, anomalies are typically identified based on a reconstruction or distance loss (Kim et al., 2022). In particular, the One-Class Deep SVDD objective (Ruff et al., 2018) is commonly used for two reasons: it can be easily combined with other neural networks, and more importantly, it generally achieves a state-of-the-art performance (Pang et al., 2021). Therefore, to detect anomalies, we train a one-class classifier by optimising the following One-Class Deep SVDD objective:
(5) $\min\limits_{\mathcal{H}}\frac{1}{M}\sum_{m=1}^{M}\lVert\mathrm{DiGCN(\mathcal{G}_{m};\mathcal{H})-\textbf{o}}\rVert_{2}^{2}+\frac{\lambda}{2}\sum_{l=1}^{L}\lVert\textbf{H}^{(l)}\rVert_{F}^{2},$
where $\textbf{H}^{(l)}$ represents the trainable parameters of DiGCN at the $l$-th layer, namely $(\Theta^{(0)(l)},\Theta^{(1)(l)},...,\Theta^{(k)(l)})^{T}$, $\mathcal{H}$ denotes $\{\textbf{H}^{(1)},...,\textbf{H}^{(L)}\}$, $\lambda>0$ represents the weight-decay hyperparameter, $\lVert\cdot\rVert_{2}$ is the Euclidean norm, and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. Moreover, o is the center of the hypersphere in the learned representation space. Ruff et al. (Ruff et al., 2018) found empirically that setting o to the average of the network representations (i.e., graph representations in our case) obtained by performing an initial forward pass is a good strategy.
Ruff et al. (Ruff et al., 2018) also pointed out, however, that One-Class Deep SVDD classification may suffer from a hypersphere collapse, which will yield trivial solutions, namely mapping all graphs to a fixed center in the representation space. To avoid a hypersphere collapse, the hypersphere center o is set to the average of the network representations, the bias terms in the neural networks are removed, and unbounded activation functions such as ReLU are preferred.
@@
(6) $score(\mathcal{G}_{m})=\lVert\mathrm{DiGCN(\mathcal{G}_{m};\mathcal{H})-\textbf{o}}\rVert_{2}.$
Table 1. Summary of datasets. #Events refers to the number of log event templates obtained using log parser Drain (He et al., 2017). #Groups means the number of generated graphs. #Anomalies represents the number of anomalous graphs. #Nodes denotes the average number of nodes in generated graphs. #Edges indicates the average number of edges in the generated graphs.
Name
#Events
#Graphs
#Anomalies
#Nodes
#Edges
HDFS
48
575,061
16,838
7
20
Hadoop
683
978
811
34
120
BGL
1848
69,251
31,374
10
30
Spirit
834
10,155
4,432
6
24
Thunderbird
1013
52,160
6,814
16
52
Table 2. Description of hyperparameters involved in OCDiGCN. Range indicates the values that we have tested, and boldfaced values represent the values suggested to use in experiments. Particularly, for the embedding dimensions: 300 is suggested for BGL and 128 for others. For the batch sizes: 32 is suggested for HDFS and 128 for others. For the training epochs: 100 for BGL and Thunderbird, 200 for HDFS, 300 for Hadoop and 500 for Spirit are suggested.
Symbol Meaning Range
$Bs$
Batch size
{16, 32, 64, 128, 256, 512, 1024, 1536, 2048, 2560}
$Op$
optimisation method
Adam, SGD
$L$
number of layers
$\{\textbf{1},2,3,4,5\}$
$\lambda$
weight decay parameter
$\{\textbf{0.0001},0.001,0.01,0.1\}$
$\eta$
learning rate
$\{0.0001,0.001,\textbf{0.01}\}$
$k$
proximity parameter
$\{\textbf{1},2\}$
$\alpha$
teleport probability
$\{0.05,\textbf{0.1},0.2\}$
$\Gamma$
fusion operation if $k\geq 2$
sum, concatenation
$Re$
readout function
mean, sum, max
$d$
embedding dimension
$\{32,64,\textbf{128},256,300\}$
$Ep$
Epochs for training
range(100,1000,50)
Training and hyperparameters: In summary, OCDiGCN is composed of an $L$-layer DiGCN architecture to learn node representations, plus a $\mathrm{Readout}(\cdot)$ function to obtain the graph representation. It is trained in an end-to-end manner via optimising the SVDD objective, which can be optimised using stochastic optimisation techniques such as Adam (Kingma and Ba, 2014). Overall, OCDiGCN takes a collection of non-anomalous graphs and a set of hyperparameters, which are outlined in Table 2, as inputs. Importantly, the pseudo-code for Logs2Graphs is given in Algorithm 1.
@@
1:Parse $D_{tr}$ and $D_{ts}$ using Drain (He et al., 2017) $\rightarrow$ Obtain parsed datasets $\hat{D}_{tr}$ and $\hat{D}_{ts}$
2:Group $\hat{D}_{tr}$ and $\hat{D}_{ts}$ based on log identifier $\rightarrow$ Obtain grouped datased $\tilde{D}_{tr}$ and $\tilde{D}_{ts}$
@@
5.3. Anomaly Explanation
Our anomaly explanation method can be regarded as a decomposition method (Yuan et al., 2022), that is, we build a score decomposition rule to distribute the prediction anomaly score to the input space. Concretely, a graph $\mathcal{G}_{m}$ is identified as anomalous if and only if its graph-level representation has a large distance to the hyper-sphere center (Equation 6). Further, the graph-level representation is obtained via a Readout()($\cdot$)( ⋅ ) function applied on the node-level representations (Equation 4). Therefore, if the Readout()($\cdot$)( ⋅ ) function is attributable (such as sum or mean), we can easily obtain the a small subset of important nodes (in the penultimate layer) whose node embeddings contribute the most to the distance. Specifically, the importance score of node $v_{j}$ (in the penultimate layer) in a graph $\mathcal{G}_{m}$ is defined as:
(7) $\frac{|score(\mathcal{G}_{m})-score(\mathcal{G}_{m}/\{\textbf{Z}_{j}\})|}{score(\mathcal{G}_{m})}$
where $score(\mathcal{G}_{m})$ is defined in Equation 6 and $score(\mathcal{G}_{m}/\{\textbf{Z}_{j}\})$ is the anomaly score by removing the embedding vector of $v_{j}$ (namely $\textbf{Z}_{j}$) when applying Readout function to obtain the graph-level representation.
Next, for each important node in the penultimate layer, we extend the LRP (Layerwise Relevance Propagation) algorithm (Bach et al., 2015) to obtain a minor set of important nodes in the input layer (this is not the contribution of our paper and we simply follow the practice in (Schwarzenberg et al., 2019; Baldassarre and Azizpour, 2019)). If certain of these nodes are connected by edges, the resulting subgraphs can provide more meaningful explanations. As the LRP method generates explanations utilizing the hidden features and model weights directly, its explanation outcomes are deemed reliable and trustworthy (Li et al., 2022).
6. Experiments

### v3

@@ 3.1 Graph-based Log Anomaly Detection
We use these definitions to define graph-based log anomaly detection:
@@
  • 1.
    A log group (here a graph) is considered a quantitative anomaly if the occurrence frequencies of some events in the group are higher or lower than expected from what is commonly observed. For example, if a file is opened (event $A$) twice, it should normally also be closed (event $B$) twice. In other words, the number of event occurrences #$\#A=\#B$ in a normal pattern and an anomaly is detected if #$\#A\neq\#B$.
  • 2.
    A log group is considered to contain sequential anomalies if the order of certain events violates the normal order pattern. For instance, a file can be closed only after it has been opened in a normal workflow. In other words, the order of event occurrences $A\rightarrow B$ is considered normal while $B\rightarrow A$ is considered anomalous.
An advantage of graph-based anomaly detection is that it can detect these two types of anomalies, but also anomalies reflected in the structures of the graphs. Moreover, no unsupervised log anomaly detection approaches represent event logs as attributed, directed, weighted graphs, which allow for even higher expressiveness than undirected graphs (and thus limiting the information loss resulting from the representation of the log files as graphs).
4 Preliminaries: Digraph Inception Convolutional Nets
To learn node representations for attributed, directed, and edge-weighted graphs, Tong et al. Tong et al. (2020) proposed Digraph Inception Convolutional Networks (DiGCN).
@@ 4 Preliminaries: Digraph Inception Convolutional Nets
$\textbf{Z}^{(k)}=\begin{cases}\textbf{X}\Theta^{(0)}&\text{$k=0$}\\ \Psi\textbf{X}\Theta^{(1)}&\text{$k=1$}\\ \Phi\textbf{X}\Theta^{(k)}&\text{$k\geq 2$},\end{cases}$ (1)
where $\Psi=\frac{1}{2}\left(\Pi^{(1)\frac{1}{2}}\textbf{P}^{(1)}\Pi^{(1)\frac{-1}{2}}+\Pi^{(1)\frac{-1}{2}}\textbf{P}^{(1)T}\Pi^{(1)\frac{1}{2}}\right)$ and $\Phi=\textbf{W}^{(k)\frac{-1}{2}}\textbf{P}^{(k)}\textbf{W}^{(k)\frac{-1}{2}}$. Particularly, $\textbf{Z}^{(k)}\in\mathcal{R}^{|\mathcal{V}|\times f}$ denotes the convolved output with $f$ output dimension, and $\Theta^{(0)},\Theta^{(1)},\Theta^{(k)}$ represent the trainable parameter matrices.
@@
$\textbf{P}^{(k)}=\begin{cases}\textbf{I}&\text{$k=0$}\\ \tilde{\textbf{D}}^{-1}\tilde{\textbf{A}}&\text{$k=1$}\\ Ins\left((\textbf{P}^{(1)})^{(k-1)}(\textbf{P}^{(1)T})^{(k-1)}\right)&\text{$k\geq 2$},\end{cases}$ (2)
@@
, with $Intersect(\cdot)$ denoting the element-wise intersection of two matrices (see Tong et al. (2020) for computation details). In addition, $\textbf{W}^{(k)}$ is the diagonalized weight matrix of $\textbf{P}^{(k)}$, and $\Pi^{(1)}$ is the approximate diagonalized eigenvector of $\textbf{P}^{(1)}$. Particularly, the approximate diagonalized eigenvector is calculated based on personalised PageRank Bahmani et al. (2010), with a parameter $\alpha$ to control the degree of conversion from a digraph to an undirected graph. We omit the details to conserve space, and refer to Tong et al. (2020) for more details.
Figure 1: The Logs2Graphs pipeline. We use attributed, directed, and weighted graphs for representing the log files with high expressiveness, and integrate representation learning and anomaly detection for accurate anomaly detection. We use off-the-shelf methods for log parsing, log grouping, and graph construction.
@@
$\textbf{Z}=\sigma\left(\Gamma\left(\textbf{Z}^{(0)},\textbf{Z}^{(1)},...,\textbf{Z}^{(k)}\right)\right),$ (3)
@@
5 Graph-Based Anomaly Detection for Event Logs
@@
First, after collecting logs from a system, the log parsing step extracts log events and log parameters from raw log messages. Since log parsing is not the primary focus of this article, we use Drain He et al. (2017) for this task. Drain is a log parsing technique with fixed depth tree, and has been shown to generally outperform its competitors Zhu et al. (2019). We make the following assumptions on the log files:
  • 1.
@@
5.1 Graph Construction
We next explain how to construct an attributed, directed, and edge-weighted graph given a group of parsed log messages, and illustrate this in Figure 2.
First, we utilise nodes to represent different log events. As a result, the number of nodes depends on the number of unique log events that occur within the log group. Second, starting from the first line of log messages in chronological order, we add a directed edge from log event $L_{i}$ to $L_{j}$ and set its edge-weight to $1$ if the next event after $L_{i}$ is $L_{j}$. If the corresponding edge already exists, we increase its edge-weight by $1$. In this manner, we obtain a labelled, directed, and edge-weighted graph.
Figure 2: The construction of an attributed, directed, and edge-weighted graph from a group of log messages.
@@ 5.1 Graph Construction
  2. 2.
    Word embedding: we use Glove Pennington et al. (2014), a pre-trained word embedding model with 200 embedding dimensions to generate a vector representation for each word in a log event;
  3. 3.
    Sentence embedding: we generate a vector representation for each log event. Since the words in a sentence are usually not of equal importance, we use Term Frequency-Inverse Document frequency (TF-IDF) Ramos et al. (2003) to measure the importance of words. As a result, the weighted sum of word embedding vectors composes the vector representation of a log event.
@@
5.2 OCDiGCN: One-Class Digraph Inception Convolutional Nets
We next describe One-Class Digraph Inception Convolutional Networks, abbreviated as OCDiGCN, a novel method for end-to-end graph-level anomaly detection. We chose to build on Digraph Inception Convolutional Networks (DiGCN) Tong et al. (2020) for their capability to handle directed graphs, which we argued previously is an advantage in graph-based log anomaly detection.
Given that DiGCN was designed for node representation learning, we repurpose it for graph representation learning as follows:
$\textbf{z}=\mathrm{Readout}(\textbf{Z}_{i}\mid i\in\{1,2,...,|\mathcal{V}|\}).$ (4)
That is, at the final iteration layer, we utilise a so-called $\mathrm{Readout}(\cdot)$ function to aggregate node vector representations to obtain a graph vector representation. Importantly, $\mathrm{Readout}(\cdot)$ can be a simple permutation-invariant function such as maximum, sum or mean, or a more advanced graph-level pooling function Ying et al. (2018).
Next, note that DiGCN work did not explicitly enable learning edge features (i.e., Y). However, as DiGCN follows the Message Passing Neural Network (MPNN) framework Gilmer et al. (2017), incorporating Y into Equation (1) and conducting computations in Equations (2-4) analogously enables learning edge features.
@@ 5.2 OCDiGCN: One-Class Digraph Inception Convolutional Nets
In graph anomaly detection, anomalies are typically identified based on a reconstruction or distance loss Kim et al. (2022). In particular, the One-Class Deep SVDD objective Ruff et al. (2018) is commonly used for two reasons: it can be easily combined with other neural networks, and more importantly, it generally achieves a state-of-the-art performance Pang et al. (2021). To detect anomalies, we thus train a one-class classifier by optimising the following One-Class Deep SVDD objective:
$\min\limits_{\mathcal{H}}\frac{1}{M}\sum_{m=1}^{M}\lVert\mathrm{DiGCN(\mathcal{G}_{m};\mathcal{H})-\textbf{o}}\rVert_{2}^{2}+\frac{\lambda}{2}\sum_{l=1}^{L}\lVert\textbf{H}^{(l)}\rVert_{F}^{2},$ (5)
where $\textbf{H}^{(l)}$ represents the trainable parameters of DiGCN at the $l$-th layer, namely $(\Theta^{(0)(l)},\Theta^{(1)(l)},...,\Theta^{(k)(l)})^{T}$, $\mathcal{H}$ denotes $\{\textbf{H}^{(1)},...,\textbf{H}^{(L)}\}$, $\lambda>0$ represents the weight-decay hyperparameter, $\lVert\cdot\rVert_{2}$ is the Euclidean norm, and $\lVert\cdot\rVert_{F}$ denotes the Frobenius norm. Moreover, o is the center of the hypersphere in the learned representation space. Ruff et al. Ruff et al. (2018) empirically found that setting o to the average of the network representations (i.e., graph representations in our case) obtained by performing an initial forward pass is a good strategy.
Ruff et al. Ruff et al. (2018) also pointed out, however, that One-Class Deep SVDD classification may suffer from a hypersphere collapse, which will yield trivial solutions, namely mapping all graphs to a fixed center in the representation space. To avoid a hypersphere collapse, the hypersphere center o is set to the average of the network representations, the bias terms in the neural networks are removed, and unbounded activation functions such as ReLU are preferred.
@@
$score(\mathcal{G}_{m})=\lVert\mathrm{DiGCN(\mathcal{G}_{m};\mathcal{H})-\textbf{o}}\rVert_{2}.$ (6)
Table 1: Summary of datasets. #Events refers to the number of log event templates obtained using log parser Drain He et al. (2017). #Groups means the number of generated graphs. #Anomalies represents the number of anomalous graphs. #Nodes denotes the average number of nodes in generated graphs. #Edges indicates the average number of edges in the generated graphs.
Name #Events #Graphs #Anomalies #Nodes #Edges HDFS 48 575,061 16,838 7 20 Hadoop 683 978 811 34 120 BGL 1848 69,251 31,374 10 30 Spirit 834 10,155 4,432 6 24 Thunderbird 1013 52,160 6,814 16 52
Table 2: Description of hyperparameters involved in OCDiGCN. Range indicates the values that we have tried on validation data, and boldfaced values are the values suggested to use in experiments. Particularly, for the embedding dimensions: 300 is suggested for BGL and 128 for others. For the batch sizes: 32 is suggested for HDFS and 128 for others. For the training epochs: 100 for BGL and Thunderbird, 200 for HDFS, 300 for Hadoop and 500 for Spirit are suggested.
Symbol Meaning Range $Bs$ Batch size {16, 32, 64, 128, 256, 512, 1024, 1536, 2048, 2560} $Op$ optimisation method Adam, SGD $L$ number of layers $\{\textbf{1},2,3,4,5\}$ $\lambda$ weight decay parameter $\{\textbf{0.0001},0.001,0.01,0.1\}$ $\eta$ learning rate $\{0.0001,0.001,\textbf{0.01}\}$ $k$ proximity parameter $\{\textbf{1},2\}$ $\alpha$ teleport probability $\{0.05,\textbf{0.1},0.2\}$ $\Gamma$ fusion operation if $k\geq 2$ sum, concatenation $Re$ readout function mean, sum, max $d$ embedding dimension $\{32,64,\textbf{128},256,300\}$ $Ep$ Epochs for training range(100,1000,50)
Training and hyperparameters: In summary, OCDiGCN is composed of an $L$-layer DiGCN architecture to learn node representations, plus a $\mathrm{Readout}(\cdot)$ function to obtain the graph representation. It is trained in an end-to-end manner via optimising the SVDD objective, which can be optimised using stochastic optimisation techniques such as Adam Kingma and Ba (2014). Overall, OCDiGCN takes a collection of non-anomalous graphs and a set of hyperparameters, which are outlined in Table 2, as inputs. The pseudo-code for Logs2Graphs is given in Algorithm 1.
@@
1:$\hat{D}_{tr},\hat{D}_{ts}\leftarrow Drain\_parse(D_{tr}),Drain\_parse(D_{ts})$
2:Group $\hat{D}_{tr}$ and $\hat{D}_{ts}$ based on log identifier $\rightarrow$ Obtain grouped dataset $\tilde{D}_{tr}$ and $\tilde{D}_{ts}$
@@
5.3 Anomaly Explanation
Our anomaly explanation method can be regarded as a decomposition method Yuan et al. (2022), that is, we build a score decomposition rule to distribute the prediction anomaly score to the input space. Concretely, a graph $\mathcal{G}_{m}$ is identified as anomalous if and only if its graph-level representation has a large distance to the hyper-sphere center (Equation 6). Further, the graph-level representation is obtained via a Readout()($\cdot$)( ⋅ ) function applied on the node-level representations (Equation 4). Therefore, if the Readout()($\cdot$)( ⋅ ) function is attributable (such as the sum or the mean), we can easily obtain the a small subset of important nodes (in the penultimate layer) whose node embeddings contribute the most to the distance. Specifically, the importance score of node $v_{j}$ (in the penultimate layer) in a graph $\mathcal{G}_{m}$ is defined as
$\frac{|score(\mathcal{G}_{m})-score(\mathcal{G}_{m}\setminus\{\textbf{Z}_{j}\})|}{score(\mathcal{G}_{m})}$ (7)
where $score(\mathcal{G}_{m})$ is defined in Equation 6 and $score(\mathcal{G}_{m}\setminus\{\textbf{Z}_{j}\})$ is the anomaly score by removing the embedding vector of $v_{j}$ (namely $\textbf{Z}_{j}$) when applying the Readout function to obtain the graph-level representation.
Next, for each important node (with a high importance score) in the penultimate layer, we extend the LRP (Layerwise Relevance Propagation) algorithm Bach et al. (2015) to obtain a minor set of important nodes in the input layer (this is not the contribution of our paper and we simply follow the practice in Schwarzenberg et al. (2019); Baldassarre and Azizpour (2019)). If certain of these nodes are connected by edges, the resulting subgraphs can provide more meaningful explanations. As the LRP method generates explanations utilizing the hidden features and model weights directly, its explanation outcomes are deemed reliable and trustworthy Li et al. (2023b).
6 Experiments